### v1

# STACKFEED: Structured Textual Actor-Critic Knowledge base editing with FEEDback

Abstract
Large Language Models (LLMs) often generate incorrect or outdated information, especially in low-resource settings or when dealing with private data. To address this, Retrieval-Augmented Generation (RAG) uses external knowledge bases (KBs), but these can also suffer from inaccuracies. We introduce STACKFEED, a novel Structured Textual Actor-Critic Knowledge base editing with FEEDback approach that iteratively refines the KB based on expert feedback using a multi-actor, centralized critic reinforcement learning framework. Each document is assigned to an actor, modeled as a ReACT agent, which performs structured edits based on document-specific targeted instructions from a centralized critic. Experimental results show that STACKFEED significantly improves KB quality and RAG system performance, enhancing accuracy by up to 8% over baselines.
1 Introduction
Large Language Models (LLMs) often produce incorrect or outdated information, particularly in low-resource settings or when handling private data. Even if the information provided is accurate, LLMs can generate hallucinated or imaginary content alongside it (Maynez et al., 2020; Zhou et al., 2021). A promising solution to address these issues is the integration of retrieval components that extract relevant information from external knowledge sources, known as Retrieval-Augmented Generation (RAG) (Chen et al., 2017; Khandelwal et al., 2020; Guu et al., 2020; Izacard et al., 2022; Shi et al., 2023). For clarity, we will refer to these external knowledge sources as Knowledge Bases (KBs). However, KBs themselves can suffer from inaccuracies, incompleteness, or outdated content. To address these challenges, there is growing interest in Knowledge Editing (KE) techniques to enhance LLMs with up-to-date and accurate knowledge.
Advancements in KE have focused on updating the model’s parameters (De Cao et al., 2021a; Meng et al., 2022; 2023), adding new parameters to model (Huang et al., 2023; Yu et al., 2024), and holding additional memory (Madaan et al., 2022; Wang et al., 2024a; b). Contrary to approaches that either update model parameters or add new parameters that require white-box access to LLMs, memory-based approaches can work with black-box access to LLMs. In similar line of thought, recently, KE approaches have also focused on refining the KBs themselves (Li et al., 2024). For example, the method proposed in Li et al. (2024) continuously updates KBs with new information, such as the current identity of the British Prime Minister. This approach demonstrates that directly editing the KB is more effective than simply adding new documents, which may coexist with outdated or inaccurate ones. Removing older documents is often not feasible, as only certain sections may be incorrect, while other parts could still provide valuable information for different queries. However, in applications like chatbots or code generation using API documentation, where updated information might not be readily available in document form, expert intervention can be crucial (Ramjee et al., 2024; Afzal et al., 2024). In such cases, expert feedback can be used to directly update the KB with accurate information when the LLM produces erroneous results.
To leverage expert or oracle feedback, we propose STACKFEED, a Structured Textual Actor-Critic Knowledge base editing with FEEDback technique. Our contributions are as follows:
  1. 1.
    Introduction of Feedback-Driven KB Editing: We present STACKFEED, a novel framework that refines the KB using structured edits based on expert feedback. This approach allows for direct, document-level updates without requiring access to LLM parameters, making it applicable to both white-box and black-box LLMs.
  2. 2.
    Multi-Actor, Centralized Critic Architecture: We design a multi-agent reinforcement learning framework where each actor is responsible for a specific document, and a centralized critic coordinates updates based on a global reward signal. This architecture ensures that document-level edits are consistent and contribute to the overall accuracy of the RAG system.
  3. 3.
    Parameterized Action Space for Document Editing: We propose a parameterized action space for each document-specific actor, enabling fine-grained control over edits, additions, and deletions within each document. This structured action space allows the actors to perform precise modifications based on expert feedback, resulting in a refined KB that better supports the RAG system.
  4. 4.
    Definition and Evaluation of KB Characteristics: We define desirable characteristics for KB refinement, including coherence, completeness, and generalizability, and introduce corresponding metrics to quantitatively assess these properties. These metrics provide a systematic way to measure the effectiveness of KB updates.
  5. 5.
    Empirical Evaluation and Performance Gains: We demonstrate that STACKFEED significantly improves the accuracy and reliability of the QA system in a variety of settings. Through extensive experiments, we show that incorporating expert feedback into document-level edits leads to a substantial reduction in error rates and enhances the KB’s ability to support accurate answer generation.
This paper is organized as follows: Section 2 reviews relevant prior work, while Section 3 presents an illustrative example to introduce and explain our approach. Section 4 details the proposed methodology, and Section 5 outlines the desired characteristics for the edited KB along with metrics for evaluation. Section 6 describes the experimental setup, and finally, Section 7 reports the results.
2 Related work
The STACKFEED framework addresses a key limitation of current RAG systems: the inability to dynamically update Knowledge Bases (KBs) without retraining or altering model parameters. Our work draws from research in Retrieval-Augmented Generation (RAG), Continual Learning, Model Editing, and feedback-driven prompt optimization, incorporating insights from Multi-Agent Reinforcement Learning (MARL) to propose an effective solution for KB editing.
Retrieval Augmented Generation (RAG): RAG systems enhance LMs by retrieving relevant knowledge from a KB based on the input query and appending it to the context, thereby addressing the limitations of standalone LMs that lack sufficient context and produce inaccurate answers (Chen et al., 2017; Khandelwal et al., 2020; Guu et al., 2020; Izacard et al., 2022; Shi et al., 2023). These systems dynamically construct contexts from unstructured KBs without modifying the LM’s internal parameters. STACKFEED further enhances RAG systems by refining the KB itself based on feedback, ensuring more accurate and up-to-date information.
Continual Learning: Continual Learning (CL) methods address the challenge of updating LMs in non-stationary environments by ensuring that new information is learned without forgetting previously acquired knowledge (Jin et al., 2022; Xu et al., 2023; Padmanabhan et al., 2023; Akyürek et al., 2024). These methods are often computationally intensive and require large-scale retraining, making them less suitable for scenarios requiring frequent updates or minimal computational resources. STACKFEED, by contrast, leverages expert feedback to perform direct edits to the KB, avoiding the need for extensive retraining.
Knowledge Editing: Knowledge Editing approaches fall into two categories: Model Editing, which modifies the LM parameters directly, and Input Editing, which updates the knowledge supplied to the model. While Model Editing efficiently alters specific facts using specialized secondary models or altering parameters (De Cao et al., 2021b; Meng et al., 2023), it struggles to ensure consistent updates across contexts (Onoe et al., 2023; Hua et al., 2024). In contrast, Input Editing modifies the KB itself, enabling updates to be reflected in outputs without changing model parameters (Madaan et al., 2022; Wang et al., 2024a; b; Li et al., 2024). STACKFEED builds on input editing techniques by leveraging expert feedback to refine the KB systematically, ensuring more accurate and consistent responses.
Prompt Optimization: With the advent of LMs, some recent works approximate gradients in text-based environments using LMs (Pryzant et al., 2023; Wang et al., 2023; Juneja et al., 2024; Gupta et al., 2024) for optimizing task prompts. STACKFEED is inspired by these approaches and generates textual reflections, similar to MetaReflection (Gupta et al., 2024) and Shinn et al. (2023), as proxies for gradients. It provides actionable guidance for document updates without the need for differentiable models. Additionally, STACKFEED adopts clustering strategies for feedback aggregation from works like UniPrompt (Juneja et al., 2024)- ensuring that actors receive coherent and non-redundant instructions.
Multi-Agent Reinforcement Learning (MARL): Multi-agent reinforcement learning (MARL) has been applied to various domains, with early research focusing on tabular methods (Busoniu et al., 2008; Canese et al., 2021; Gronauer & Diepold, 2022) and later expanding to deep learning techniques for high-dimensional inputs (Tampuu et al., 2017; Leibo et al., 2017). Studies have explored independent Q-learning (Tan, 1993), agent communication (Foerster et al., 2016; Das et al., 2017), and centralized training with decentralized execution (Gupta et al., 2017). However, most of these approaches do not address the critical challenge of multi-agent credit assignment. Actor-critic methods have been introduced to overcome this limitation by employing centralized critics with decentralized actors  (Foerster et al., 2018; Iqbal & Sha, 2019; Wang et al., 2021; Chen et al., 2023). STACKFEED extends such actor-critic framework to operate directly on textual content, using the centralized critic to decompose feedback into actionable textual gradients for each document-specific actor.
In the next section, we provide an example to illustrate the KB editing problem, while also providing an overview of STACKFEED.
3 Example and Overview
Figure 1 illustrates our technique applied to the ARKS Pony domain (Su et al., 2024a), where a knowledge base (KB) for the low-resource programming language Pony supports a natural language-to-code task. Due to Pony’s rarity, language models often generate code that fails to compile. To address this, we use the Pony compiler as an expert to provide feedback in the form of compile errors.
1
Evaluating the Knowledge Base State: We start with an initial KB, including documents like builtin-array.md. The system retrieves relevant documents based on the given task (e.g., counting non-inversions in an array) and generates a program, which is evaluated by the compiler, resulting in feedback (e.g., compile errors).
2
Centralized Feedback Analysis: We analyze compile errors to generate reflections that explain why the errors occurred. For instance, if the apply method in the Array class is partial and may raise an error, the reflection suggests adding a ? to handle potential failures. These reflections are matched to the documents they pertain to, refining the understanding of errors.
3
Distributing Gradients: Reflections are generalized into gradients, which summarize modifications needed for each document. For example, the theme might be the partial nature of functions like apply and update, which need better error handling in the documentation.
4
Generating Edit Actions: Gradients are converted into structured edit actions, such as adding or modifying content in specific sections of the documents.
5
Re-evaluation and MCTS Search: After edits are applied, the KB is re-evaluated, generating new feedback and a reward score. This score guides a Monte Carlo Tree Search (MCTS) to explore different states of the KB, iterating through steps 1- 4 to progressively refine the KB and improve the system’s overall performance.
Task: Given an array of integers nums, write a function that returns the number of good pairs. A pair (i, j) is called good if nums[i] == nums[j] and i < j. Retrieved Documents:
    builtin-array.md
    collections-persistent-vec.md
    math-fibonacci.md
    random-dice.md
Output Program:
fun numIdenticalPairs(ns: Array[I32]): I32 =>
  var count: I32 = 0
  for i in Range(0, ns.size() - 1) do
    for j in Range(i + 1, ns.size()) do
      if ns(i) == ns(j) then
        count = count + 1
Simulate
Task
Knowledge Base
    builtin-array.md
    collections-persistent-vec.md
    math-fibonacci.md
Retrieval Expert (Compiler) Feedback
main.pony:15:16: call is not partial but the method is
  - a question mark is required after this call
      if nums(i) == nums(j) then
main.pony:15:27: call is not partial but the method is
  - a question mark is required after this call
      if nums(i) == nums(j) then
main.pony:11:42: function signature is not marked as
  partial but the function body can raise an error
      fun numIdenticalPairs(nums: Array[I32]): I32 =>
Feedback Reflection: The output code contains the following errors: 1. The array access calls nums(i) and nums(j) are not marked as partial, but they should be as the apply function in the Array class is partial and can throw errors. The corrected code is: nums(i)? = nums(j)? To further illustrate the fix, consider the following simplified example: Code with error: fun get(arr: Array[I32], index: USize): I32 ? => arr(index) Fixed code: fun get(arr: Array[I32], index: USize): I32 ? => arr(index)? ... Selection: builtin-array.md
Reasoning: This file contains information about the Array class, including apply...
Collected selection reasoning for builtin-array.md Reflection and Selection Document-wise Partial Gradient: builtin-array.md is a KB document detailing the use of the methods and functions of the Array class...
Method Availability and Error Handling: The current file lacks explicit guidance on the partiality of methods and the use of ? operator. The omission can lead developers to... The file should accurately reflect the partiality of methods like apply and update...
Aggregate Document Gradients Edit: Add Section
When working with arrays in Pony, it is important to understand that some methods are partial, meaning the may not succeed in all cases and might raise an error. To safely handle these errors, Pony provides the ? operator...
The apply method on arrays is partial and will raise an error in the index is out of bounds...
For example, the ‘apply‘ method on arrays is partial and requires error handling:
‘‘‘pony try my_array.apply(42) == 7 end ‘‘‘ In this snippet, if the index 42 is out of bounds the ‘apply‘ method will raise an error, which is then handled by the ‘try‘ block, preventing the program from crashing.
Generate Edits
Figure 1: Example of the STACKFEED in the ARKS Pony scenario
4 Methodology
We will start by describing a typical Retrieval-Augmented Generation (RAG) system over unstructured Knowledge Bases.
Errors in such systems can arise from multiple components: 1) the LLM $\mathcal{B}$ might fail to reason correctly over the provided information, 2) the retriever $R$ might not select the right set of relevant documents from $\mathcal{K}$, or 3) the knowledge base $\mathcal{K}$ itself might contain incorrect or incomplete information. We assume an expert is monitoring the system, identifying when answers are incorrect, determining which component is at fault, and providing feedback on why the answer is incorrect and what the correct answer must be.
This work focuses on scenarios where incorrect answers result from issues in the Knowledge Base ($\mathcal{K}$). Our goal is to improve $\mathcal{K}$ by addressing mistakes in $\mathcal{K}$ and filling in missing information based on expert feedback, thus enhancing the RAG system’s performance on future queries.
4.1 Problem Formulation
We are provided with a training set $T=\{(q_{i},o_{i},c_{i},f_{i})\}_{i=1}^{l}$, where $q_{i}$ is a user query, $o_{i}$ is the RAG system’s answer, $c_{i}$ is the correct answer, and $f_{i}$ is an optional expert feedback on incorrect answers. We also assume access to a scoring function $g$, which compare $o_{i}$ and $c_{i}$ to output a score. The objective is to optimize the knowledge base $\mathcal{K}$ to maximize the sum of the scores for all queries in the training set:
$\mathcal{K}^{*}=\arg\max_{\mathcal{K}}\frac{1}{|T|}\sum_{(q_{i},a_{i},c_{i},f_{i})\in T}g(\mathcal{B}(q_{i},\Gamma(q_{i},\mathcal{K})),c_{i})$ (1)
In the next section, we show how such an objective can be seen as a state search problem.
4.2 Knowledge Base Editing as State Search
In our problem setting, the Knowledge Base ($\mathcal{K}$) is defined as a collection of documents $\mathcal{K}=\{D_{i}\}_{i=1}^{n}$. We assume each document consists of a number of chunks of text and can be represented as $D_{i}=[c_{ij}]$. The state $s\in\mathcal{S}$ of the system is represented by the current configuration of the KB, i.e., the content of all documents in $\mathcal{K}$.
Given a query $q_{i}$ and a set of retrieved documents $\Gamma(q_{i},\mathcal{K})$, the LLM $\mathcal{B}$ generates an answer $o_{i}$. When errors arise due to incomplete or incorrect information in the retrieved documents, our goal is to identify the optimal configuration of $\mathcal{K}$ that improves the accuracy of the system’s responses. Thus, we define our state search problem as finding the best state $s^{*}$ of the KB.
State Space: The state space $\mathcal{S}$ encompasses all possible configurations of the KB. Each state $s$ corresponds to a particular set of document contents, represented as: $s=\{D_{i}\}_{i=1}^{n},$ where $D_{i}$ denotes the content of document $i$ and $n$ is the number of documents in $\mathcal{K}$. The state $s$ captures the overall structure and content of the KB at any given point. We set $s_{0}=\mathcal{K}$.
State Transition Function: The state transition function $\mathcal{T}(s,u)$ defines how the KB changes in response to the action $u$ taken by the agent. Each action contains modifications to one or more documents within the KB, resulting in a new KB configuration. The state transition is formalized as: $s^{\prime}=\mathcal{T}(s,u),$ where $s^{\prime}$ is the new state of the KB after applying $u$.
Action Space: The action space $\mathcal{A}$ consists of list of diffs $d_{i}$ corresponding to each document $D_{i}$. Essentially, $u=[d_{i}]_{i=1}^{|\mathcal{K}|}$.
Environment: We model the environment simply as a ‘‘patch’’ function, that takes the diff generated by the agent and patches the KB to produce the new state.
Optimization Objective: Following Equation 1, our objective then is to find the optimal state $s^{*}$ of the KB that maximizes the overall performance of the RAG system, as measured by a global reward function $R$. The optimization problem is formulated as:
$s^{*}=\arg\max_{s\in\mathcal{S}}R(s)=\arg\max_{s\in\mathcal{S}}\frac{1}{|T|}\sum_{(q_{i},a_{i},c_{i},f_{i})\in T}g(\mathcal{B}(q_{i},\Gamma(q_{i},s)),c_{i})$ (2)
where $R(s)$ represents the cumulative reward of the KB state $s$, reflecting its ability to support accurate and complete responses for a set of queries.
The reward function $R(s)$ is derived from the expert feedback on the system’s generated answers and captures improvements in terms of correctness, coherence, and completeness of the information in the KB. By optimizing for $s^{*}$, we ensure that the final state of the KB maximizes the overall accuracy and effectiveness of the RAG system, rather than focusing on an intermediate sequence of state transitions.
In summary, the state search formulation defines the problem of finding the optimal state $s^{*}$ of the KB that maximizes the system’s performance. This approach enables us to make targeted, feedback-driven edits to the KB and achieve a refined, high-quality knowledge base that better supports accurate answer generation.
Monte Carlo Tree Search: We employ Monte Carlo Tree Search (MCTS) similar to PromptAgent  (Wang et al., 2023) to search for the optimal state $s^{*}$. However, this introduces several challenges: (1) The search space for all possible KB edits is vastly larger than that of standard prompt edits typically explored in the literature (Pryzant et al., 2023; Wang et al., 2023; Juneja et al., 2024; Gupta et al., 2024), making exhaustive search infeasible. (2) Generating actions and subsequent states, as done in methods like PromptAgent , is difficult in the KB editing context since fitting the entire KB into the prompt of a language model is impractical. Despite advancements in handling long contexts (Wang et al., 2020; Kitaev et al., 2020; Press et al., 2022; Su et al., 2024b), these models often struggle to leverage extensive contexts effectively Liu et al. (2024). (3) Finally, the LM would need to output the entire edited KB, which is challenging due to the inherent difficulty LMs face in generating long, coherent outputs (Bai et al., 2024).
To address these challenges, we decouple the KB edits by isolating document-level modifications based on the required updates. Since individual documents can be large, we further break down the edits into manageable sections, enabling a structured editing mechanism that focuses on specific portions of a document at a time. In the next section, we introduce STACKFEED, an agent designed to efficiently perform these structured edits based on feedback.
4.3 STACKFEED
Figure 2: STACKFEED  Multi-actor, centralized critic architecture: On receiving a reward from the environment, the critic generates a reflection over the failures to calculate the textual gradient $\nabla$. The critic uses this reflection to select the documents responsible for the error and proceeds to assigns credit to the actors in the form of document-wise reflections. The actors then proceed to iteratively edit the documents. All the document-wise edits are then pooled to define the KB edit.
The proposed approach STACKFEED is designed to enhance a RAG system by refining the underlying Knowledge Base ($\mathcal{K}$) using expert feedback. Our approach employs a multi-actor, centralized critic architecture, where each actor is responsible for making updates to a specific document within $\mathcal{K}$, and a centralized critic uses global feedback to coordinate these updates. The objective is to iteratively improve $\mathcal{K}$ such that the overall accuracy of the RAG system is maximized.
4.3.1 Reward Signal
For a given query $q_{i}$ and the generated answer $o_{i}$, the expert provides feedback $(c_{i},f_{i})$ that includesa ground truth answer $c_{i}$ and qualitative expert feedback $f_{i}$ on any errors. The global reward signal is derived from $c_{i}$ as per the scoring function $s$ (Refer Equation 2).
4.3.2 KB Editing Agent
To effectively incorporate expert feedback, we employ a multi-actor, centralized critic architecture.
Centralized Critic: The centralized critic, denoted as $C$, is responsible for evaluating the overall performance of the RAG system based on the global reward signal $r$ derived from expert feedback. The critic analyzes the feedback received given the current state $s$ of $\mathcal{K}$. The critic’s analysis is then used to provide tailored reflections to each actor, guiding document updates.
The centralized critic aggregates the reward signal across multiple queries to generate a holistic evaluation of $\mathcal{K}$.
$R(s)=\frac{1}{|T|}\sum_{(q_{i},a_{i},c_{i},f_{i})\in T}g(\mathcal{B}(q_{i},\Gamma(q_{i},s)),c_{i})$ (3)
To generate feedback for the documents, the critic needs to take gradient of this reward with respect to the documents. This would give us,
$\partial_{j}=\frac{\partial R(s)}{\partial D_{j}}=\frac{1}{|T|}\sum_{(q_{i},a_{i},c_{i},f_{i})\in T}\frac{\partial}{\partial D_{j}}g(\mathcal{B}(q_{i},\Gamma(q_{i},s)),c_{i})$ (4)
Figure 2 illustrates the environmental interaction of the actor-critic model. Following methodologies in prior works (Pryzant et al., 2023; Juneja et al., 2024; Gupta et al., 2024), we use LLMs to generate an overall text gradient $\nabla$ over each failing example. The critic first identifies and select which documents in $\Gamma(q_{i},s)$ are responsible for any inaccuracies in $o_{i}$. Reflections are then generated for these documents based on the correct answer, expert feedback and the text gradient. However, as shown in Equation 4, we need to aggregate these reflections across all queries. Instead of a simple concatenation, we adopt the clustering approach similar to Juneja et al. (2024), producing generalized reflections that effectively capture the core insights from multiple queries. These aggregated reflections can be effectively considered as the partial textual gradient $\partial$ with respect to the document. These partial gradients are provided as feedback to the document-specific actor $A_{j}$, which then perform the actions to edit the specific documents.
Actors: Each document $D_{i}\in\mathcal{K}$ is managed by a distinct actor, $A_{i}$, which is modeled as a ReACT agent Yao et al. (2023) responsible for making structured edits to its document. Each actor operates independently, receiving reflections from the centralized critic on how to modify the content of $D_{i}=[c_{ij}]$.The actors need to only update these chunks as needed. The set of possible actions includes:
  • $\mathsf{EditChunk}$: The action is defined as $\mathsf{EditChunk}(j,t_{j})$, where $j$ indicates which chunk $c_{ij}$ of $D_{i}$ to modify, and $t_{j}$ is the updated content for the chunk.
  • $\mathsf{AddChunk}$: The action is defined as $\mathsf{AddChunk}(n_{j},t_{j})$, where $n_{j}$ indicates the name of the new chunk, and $t_{j}$ is the content for the chunk.
  • $\mathsf{DeleteChunk}$: The action is defined as $\mathsf{DeleteChunk}(j)$, where $j$ specifies which chunk $c_{ij}$ of $D_{i}$ to remove.
This parameterized action space allows the actors to perform precise edits within the document, ensuring that the refinement process is both flexible and context-specific. Each actor leverages its local state $s_{i}$ and the document-specific feedback from the critic to produce a sequence of structured edits, ensuring that modifications are consistent and contribute towards enhancing the document’s relevance and completeness.
The ReACT agent utilizes these reflections and iteratively generates a trajectory $t_{0}={a_{0},a_{1},a_{2}\cdot a_{n}}$ of edit actions to the document until the errors are resolved or the knowledge gaps are filled. This controlled editing process improves the accuracy of the RAG system by ensuring that the KB contains up-to-date and relevant information. After the completion of the actor runs, we generate the edit diffs for each document $d_{i}$ and pool them to generate the KB edit action $u=[d_{i}]_{i=1}^{|\mathcal{K}|}$
However, there might be many ways to edit a KB and we may need to have some desirable characteristics for the edited KB. In the next section, we discuss what those desirable characteristics could be and how we might measure them.
5 Evaluating Knowledge Base Editing Quality
A Knowledge Base should be complete with respect to a task - it should contain all the information necessary to assist the RAG system to solve the task at hand. Given the open-ended nature of tasks that typical RAG agents are designed for, it is hard to quantify a closed-form metric of completeness. That said, an ideal Knowledge Base editing system should at least be able to incorporate as much external feedback as possible.
Further, It will be extremely undesirable for any Knowledge Base to only help the RAG system for a small subset of tasks. Given the tendencies for data-driven techniques to over-fit on the train-set distribution, it is important that knowledge base edits are generalizable to unseen examples.
Lastly, given the semantic and textual nature of the Knowledge Base, it is important that the documents in the Knowledge base are coherent and consistent throughout. This not only makes the document interpretable for human consumption, it also help reduce in-context noise during LLM inteference, which has been shown to affect LLM performance (Liu et al., 2024).
6 Experimental Setup
6.1 Baseline
While there has been a rich body of works in the area of knowledge editing and prompt optimization, to the best of our knowledge, STACKFEED is the first work targeting the feedback-driven textual Knowledge Base Editing problem. Therefore, to perform a holistic evaluation of STACKFEED we implement - PromptAgent-E, an extension of PromptAgent   Wang et al. (2023) for the KB editing task. PromptAgent  formulates prompt optimization as a strategic planning problem using Monte Carlo Tree Search (MCTS). At a high-level our baseline approach, PromptAgent -E creates separate PromptAgent -style agents to optimize specific document in the KB. To minimize spurious edits in the Knowledge Base, we restrict PromptAgent -E to only optimize documents that were part of the retrievals for more than 2 training sample. After identifying the best nodes for each of the document-wise runs, we put them back in the knowledge base to generate the new version of the KB. In contrast to STACKFEED, PromptAgent -E can be seen as a collection of document-wise Independent Actor-Critic models  (Foerster et al., 2017). We present in-depth comparisons between PromptAgent -E and STACKFEED in Section 7
6.2 Datasets
Knowledge Base Editing can be useful for scenarios where the KB is 1. Incomplete, or 2. Incorrect . We evaluate STACKFEED  on 5 datasets spanning these different settings.
6.2.1 Incomplete Knowledge Base
Dataset Train Eval Test Documents
Pony 31 32 45 601
Ring 26 27 39 577
ScipyM 22 22 98 3921
TensorflowM 9 9 26 5859
CLARKS News 30 30 60 138
Table 1: Dataset Statistics
We adapt two code generation datasets from ARKS (Su et al., 2024a), namely ARKS-Pony and ARKS-Ring. The dataset consists of LeetCode problems and their solutions in low-resource languages Pony and Ring respectively. Each datapoint is supplemented with a corresponding language documentation, with execution accuracy as the success metric and execution failures as feedback to the system. Given that these language don’t appear prominently in LLM pre-training data, the performance of code generation RAG agents on these datasets depends significantly on the quality of the Knowledge Base. However, given that these languages have smaller communities, their documentation isn’t as well maintained and often lack critical information. . For the purpose of evaluation on these datasets, we split them into train, eval, test splits as specified in Table  1. To ensure that we have a good representation of failure cases during training, we first execute the RAG pipeline on the entire dataset and divide the failures at random in a 1:1:2 ratio for train, eval and test respectively. All the datapoints with successful execution match are put in the test split. We use the compiler feedback from the executions as the expert feedback to the STACKFEED system.
6.2.2 Incorrect Knowledge Base
For evaluating under this setting, we leverage the ARKS-ScipyM and ARKS-TensorflowM datasets from ARKS and the CLARK-news dataset from Erase (Li et al., 2024). The ARKS datasets consist of data science problems sourced from the DS-1000 dataset  (Lai et al., 2022), which are to be solved by artificially perturbed versions of scipy and tensorflow libraries respectively, while referring to the original unperturbed documentation. Similar to Pony and Ring, we use the execution accuracy on a test bench as a success metric and use compiler outcome as expert feedback. We also follow a similar approach for data splitting.
Dataset Ring Pony SciPy Tensorflow CLARK-news
Acc $\sigma$ Acc $\sigma$ Acc $\sigma$ Acc $\sigma$ Acc $\sigma$
Base KB 30.77 2.09 29.99 1.57 52.04 0.00 28.88 2.18 26.27 1.20
PromptAgent-E 33.33 2.81 32.22 1.57 53.40 3.12 47.77 3.57 28.80 2.39
STACKFEED 36.75 1.21 37.04 1.28 59.38 1.22 53.84 3.11 37.28 1.69
Table 2: Generalization performance comparison between STACKFEED and baseline models across multiple datasets, reported as accuracy percentages (higher is better).
While fact retrieval is one of the most popular use cases of RAG systems, evolving nature of information requires us to keep the knowledge bases up to date. To simulate these dynamic factual knowledge updates we use the CLARKS-news dataset from Erase (Li et al., 2024) which contains questions and their respective answers extracted from Wikidata at different timestamps. Each timestamp is characterized by a set of articles that were added in the data at that time. For our evaluation, we pool all the questions whose answers changed for the first time at a given timestamp and split them across train, eval and test splits in a 1:1:2 ratio (Table 1).
6.3 Evaluation Metrics
In section 5 we discussed the desirable properties of a Knowledge Base edit. We leverage these properties to design 3 metrics for the KB Editing problem as follows:
Completeness: We use the train set accuracy to estimate the degree of expert feedback incorporated in the learnt Knowledge Base.
Generalization: To estimate the degree of generalization of our Knowledge Base edits, we use the held out test set accuracy.
Coherence: To quantify the degree of coherence of the KB, we first calculate a document-wise coherence score using G-Eval (Liu et al., 2023) with GPT4-1106-preview  as the judge model. The G-eval prompt assigns a 1-5 score to the diff of changes with respect to the original document, checking for thematic similarity of the diff. We pool all the edited documents for a KB edit and average there respective coherence score to define the KB coherence metric.
6.4 System Configurations
MCTS parameters: We use the Upper Confidence bounds applied to Trees (UCT) algorithm for selecting expansion nodes, enabling effective exploration and exploitation of the KB state space. For our experiments, we set a maximum search depth of $3$, an expansion width of $3$, and a maximum of $5$ iterations. The UCT exploration constant is set to $2.5$. These parameters were chosen to balance the computational cost and the need for adequate exploration. A depth of $3$ ensures that the search can explore sufficient variations in the KB states without unnecessary expansion, while an expansion width of $3$ allows a moderate number of candidate states to be evaluated at each step. Similarly, $5$ iterations provide enough opportunity to refine the state search, and the UCT constant of $2.5$ encourages sufficient exploration in early stages while converging towards high-reward states in later stages. For unstructured data, the documents are chunked after every 50 lines and then edit the chunks.
RAG System: For the purpose of our evaluations, we setup a generic RAG system which uses an embedding similarity for semantic retrieval. Additionally, in lines with prior works like  (Zhang et al., 2023) for coding related tasks, we use an iterative retrieval setup wherein we first generate a code using naive retrieval and then query the database again with both the question and generated code to improve the quality of retrieval before generating the final result.
LLM configs: We use OpenAI-Text-Embedding-3-Large  as the embedding model with dimensions size of 3072 and use cosine similarity as a metric of embedding match for ranking. To account for the 8191 max input limit, we create document chunks of at most 7500 tokens. For the reasoning model, we use GPT4-1106-preview, with a temperature of 00. Since LLMs are known to perform poorly with longer context input (Liu et al., 2024), we restrict the max token budget for retrievals at 18000 tokens and remove any lower ranked retrieval to fit this token budget.
7 Results
Completeness( in %) Coherence (1-5, higher is better)
Dataset Ring Pony SciPy Tensorflow CLARK-news Ring Pony SciPy Tensorflow CLARK-news
PromptAgent-E 4.27 3.22 33.33 33.33 11.86 4.33 1.86 2.0 4.0 1
STACKFEED 8.98 9.68 31.38 44.44 13.79 4.67 4.6 4.30 4.0 1
Table 3: Completeness and coherence comparison between STACKFEED and baseline models across multiple datasets. Completeness is reported as accuracy percentages (higher is better), while coherence is measured on a scale of 1-5 (higher is better).
7.1 Completeness and Generalization
We observe consistent improvements over the PromptAgent-E baseline in completeness and generalizability scores, with STACKFEED achieving approximately 2x performance gains on Ring and Pony datasets. However, feedback incorporation remains limited, likely due to suboptimal retrieval or limited document-query associations hindering learnability of edit. STACKFEED also demonstrates higher generalizability and lower variance, attributed to its structured and focused document edits that enhance coherence.
7.2 STACKFEED makes high quality coherent edits
As seen in Table 3, STACKFEED produces edits with a coherence score of 4 or higher for most datasets. For KBs which need long term maintenance (like language and code documentation as seen in the ARKS datasets), STACKFEED makes more coherent edits compared to the baseline. This is especially true for long documents as seen in the ARKS Pony dataset. For news-article like dataset like CLARK-news with factual edits. Incoherency is naturally induced when the facts of the article are changed. For instance, an article on the coronation of a king will lose coherency when the article is updated to add information about the coronation of a new king.
8 Conclusion
We introduced STACKFEED, a novel framework for refining Knowledge Bases (KBs) in Retrieval-Augmented Generation (RAG) systems using a multi-actor, centralized critic architecture. STACKFEED enables efficient KB updates without retraining or altering model parameters by leveraging feedback-driven structured edits and textual gradients.
Our approach achieved superior performance in preserving knowledge base (KB) coherence, consistency, and completeness, resulting in enhanced RAG system responses. Nonetheless, there remains considerable potential for further advancements. Future work will focus on refining these three metrics to elevate system performance even further.
Appendix A Appendix
A.1 Prompts used in STACKFEED
[…]
""" There exists a Language Model based software named CodeRAG that automatically does the following task for a developer: task - task_desc
CodeRAG uses a knowledge base to perform this task: kb_desc
A developer used CodeRAG to perform the task on multiple files, and CodeRAG made some errors on them.
Here is one knowledge base file that was involved in these errors: """ for i, file in enumerate(kb_files): prompt += f""" File i+1: id: file[’id’] content:
n<file>
nfile[’content’]
n</file>
n""" if "special_notes" in file and file["special_notes"] != "": prompt += f"""
nspecial_notes: file[’special_notes’]"""
""" The following are the reflections on the errors made by CodeRAG: reflections_str
The reflections show the relationship of the file with the errors made by CodeRAG. If the file is named "None," it means the information about the error on which the reflection is based does not fully fit any knowledge base file.
Your task is to use the reflections on the errors made by CodeRAG and provide a generalization on the issues with the file and how it can be improved to prevent the errors.
You should mention common issues found in the reflections and provide a plan for improving the knowledge base files to prevent future errors. Use the reflections to suggest additions or changes in the file, explaining what new content should be added to prevent errors. Before suggesting your plan, give context on the errors using code snippets and other relevant information from the reflections.
You have a scratchpad to reason and plan your generalization. Your scratchpad is for your use only and will not be shared with anyone else. The scratchpad is represented by the <scratchpad></scratchpad> tags.
Your generalization should follow this format: <scratchpad> The contents of the scratchpad </scratchpad> <generalization> Your generalization for this file </generalization>
You must provide the filled-out scratchpad and generalization in the above format.
General guidelines: 1. Carefully analyze the reflections to understand the errors CodeRAG is making. 2. "None" is a special file, representing that to fix the error, the information should be in a new file. """
Figure 3: Generalization Stage Prompt
[…]
""" Your task is to reflect upon the errors made by CodeRAG based on the user feedback and provide a reflection on the role of the knowledge base files in the making of those errors.
Your reflection should be very specific to the knowledge base files as these reflections will be used to improve the knowledge base files to prevent such errors in the future.
There may be other causes for the error, but you should only focus on whether the knowledge base files could have prevented the error.
You should also provide a way for improving the knowledge base files to prevent the error from happening again.
You should try and see if there is any error in the information provided by the knowledge base or if the knowledge base is missing some information that could have prevented the error.
You also have to figure out if the file should be edited or not. That you do through the needs_editing flag.
You have a scratchpad in which you can reason and plan your reflection. Your scratchpad is for your use only and will not be shared with anyone else. This scratchpad is represented by the <scratchpad> tags.
Your output should be in the following format:
<scratchpad>
The contents of the scratchpad
</scratchpad>
<reflection>
<File 1>
File: Name of the first file
needs_editing: True/False
Reflection: The reflection for this file
</File 1>
<File 2>
File: Name of the second file
needs_editing: True/False
Reflection: The reflection for this file
</File 2>
</reflection>
You have to provide the filled-out scratchpad and the reflection in the above-described formats. You have to reflect on all the files that were extracted for the code file.
Here are some general guidelines to follow:
  1. 1.
    You should first analyze the question, the test bench, the feedback, and the output to understand the error made by CodeRAG.
  2. 2.
    Then you should carefully analyze the knowledge base files to see if the theme and the contents of any knowledge base file are relevant to the error. Particularly, you should look out for files that have a factual error related to the error or are missing some information which should have been in the file according to the theme of the file.
    1. (a)
      Read the content of the file and understand the theme of the file. The theme of this file is of course based on the file ID and the content of the file but you should also consider its positioning in the knowledge base. That means you should consider the other files that were extracted for the code file and see how this file fits in with them. For example, if the file is a very basic general guide to the task with other files providing more detailed information, then it would make sense for this file to not have detailed information about specific cases.
    2. (b)
      See if the file has any information related to the error. Check for relevant keywords and how the file might have biased the language model to make the error.
    3. (c)
      If the file has information related to the error, see if the information is correct and complete. If the information is incorrect or incomplete, the file is responsible for the error.
    4. (d)
      If it doesnt́ have information related to the error, check if it makes sense for the file to have information related to the error. If it doesn’t make sense, the file is not responsible for the error. When deciding this, check whether the information would be better suited in any of the other knowledge base files. If the missing info fits better in another file, then deem this file to not be responsible for the error as the missing content can be better placed in the other file.
    5. (e)
      If the file is responsible for the error, explain the error in your reflection and set the needs_editing flag to True. And if the file is not responsible for the error, set the needs_editing flag to False.
  3. 3.
    If none of the files have any error or if you think the content for the error should be in a new file, put a file with the name ‘‘None’’ in your reflection and for its reflection, describe the error and mention why it is not due to the knowledge base files. For the ‘‘None’’ file, the needs_editing flag should always be set to True. The ‘‘None’’ file should be placed as File n+1 where n is the number of files extracted for the code file.
  4. 4.
    Choose the least number of files for editing, we want to change as few files as we can for any error. For example, if we have 5 knowledge base files, unless very extreme cases, we wouldn’t want to set the needs_editing flag as True on more than 2 files. Figure out what the most relevant files for the error are and focus on them.
  5. 5.
    When you choose to edit multiple files, you should make sure that their involvements in the error are distinct and not overlapping. If they are overlapping, think about whether changing one file would be enough to fix the error.
Figure 4: Selection Stage Prompt
[…]
""" There exists a Language Model based software named CodeRAG that automatically does the following task for a developer:
{test_bench_code}
The test bench code gives a code where a function must be inserted and then it is tested with some
test cases.
CodeRAG then outputted the following code to answer the question:
if task_desc != "":
    prompt += f"""
{task} - {task_desc}
else:
    prompt += f"""
{task}
prompt += f"""
The developer used CodeRAG for a question. The question is as follows:
{query}
In the question, the developer provided the following test bench code:
{test_bench_code}
The test bench code gives a code where a function must be inserted and then it is
tested with some test cases.
CodeRAG then outputted the following code to answer the question:
{output_code}
Based on the above output, the developer gave the following feedback to CodeRAG:
{feedback}
CodeRAG uses a knowledge base to do this task
{kb_desc}
The following files were extracted for this particular code file (the content of
each file is surrounded in <file></file> tags):
for i, instruction in enumerate(instructions):
    prompt += f"""
File {i+1}:
id: "{instruction[’id’]}"
[…]
    if "special_notes" in instruction and instruction["special_notes"] != "":
[…]
prompt += """
Your task is to reflect upon the errors made by CodeRAG based on the user feedback.
You have to explain in detail the error made by CodeRAG. The reflection should be
very specific to the question, the output code and the feedback.
You should start by explaining the question that CodeRAG was asked to solve before talking about the error.
Your reflection should have relevant code snippets from the output
code which have errors and what should be done to fix them.
You should also add a small code example to demonstrate the error and potential methods to fix it.
You can talk about multiple different methods here to address the error.
You have a scratchpad in which you can reason and plan your reflection.
Your scratchpad is for your use only and will not be shared with anyone else.
Your reflection should be in the following format:
<scratchpad>
The contents of the scratchpad
</scratchpad>
<reflection>
Your reflection
</reflection>
Figure 5: Reflection Stage Prompt
[…]
There exists a Language Model based software named CodeRAG that automatically does the following task
for a developer:
{task} - {task_desc}
CodeRAG was used to perform the task on a specific repository.
The following files in the repository were edited by CodeRAG:
<files>
{files}
</files>
A developer has provided the following feedback on CodeRAG’s output:
<feedback>
{feedback}
</feedback>
Your task is to parse this feedback.
You must separate the larger feedback into smaller feedbacks, where each feedback corresponds
to a specific file in the repository.
**DO NOT** change the content of the feedback. Your job is only to split the feedback into file-level
feedbacks, without altering the feedback’s content in any way.
You should respond in the following output format:
‘‘‘json
    "task": The task name, like {task},
    "num_files": The number of files mentioned in the feedback,
    "feedback_files": [
            "target_file": The file name that the feedback is about, it should be the
            whole path of the file, leave it empty if you are not sure which file the feedback is about. If
            the file name is not empty,
            it should be one of the files in the repository.
            If the file name does not exist in the repository, put an empty string here,
            "num_feedbacks": The number of feedbacks for this file,
            "feedbacks": [
                    "feedback_tone": The tone of the feedback, can take the values ‘positive‘ or
                    ‘negative‘,
                    "target_spans": The spans of the code that the feedback is about.
                    It is a list of json objects with two keys, ’start’ and ’end’.
                    Put an empty list if you are not sure which code span the feedback is about,
                    "feedback": Description of the feedback,
Make sure to include the backticks (‘‘‘) surrounding the output. They are needed for further parsing of your output. """
Figure 6: Feedback Parsing Prompt